OpenStudy (anonymous):

HELP! NEED HELP FAST! The sets of numbers 6, 8, 10 and 5, 12, 13 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (anonymous):

Well, think of how the Pythagorean Theorem utilizes an a and b to obtain c and work backwards. \[a^2 + b^2 = c^2\] \[6^2 + 8^2 = 10^2\] \[36 + 64 = 100\] \[100 = c^2\]

OpenStudy (anonymous):

i still dont understand how to explain it

OpenStudy (anonymous):

@bakus522

OpenStudy (anonymous):

the a, b, and c of a triangle make a triple. If you can prove that these are the a, b, and c of a triangle, then they are triples.

OpenStudy (anonymous):

Okay thank You!

OpenStudy (anonymous):

6, 8, and 10 are all lengths of a triangle. 6 units- the left side 8 units- the right side 10 units- the bottom a² + b² = c² 6² + 8² = 10² 36 + 64 + 100? YES. If the equation is correct, it is a Pythagorean triple. || IS THAT GOOD? @bakus522 ||

OpenStudy (anonymous):

Yeah! That's pretty much it.

OpenStudy (anonymous):

alright! Thanks so much!

OpenStudy (anonymous):

No problem.